# Graph atomic cluster expansion for foundational machine learning interatomic potentials

### v1

@@ I Introduction
The ability to predict materials properties from atomistic simulations is essential for modern materials design. Machine learning interatomic potentials (MLIPs), trained on data from electronic structure methods like density functional theory, have recently emerged as a powerful tool, achieving excellent accuracy for diverse systems Behler and Parrinello (2007); Bartók et al. (2010); Shapeev (2016); Drautz (2019); Unke et al. (2021); Musaelian et al. (2023); Jacobs et al. (2025); Kulichenko et al. (2024); Wang et al. (2024); Thiemann et al. (2024); Yuan et al. (2025). However, the applicability of conventional MLIPs is constrained by their limited elemental scope. Extending an existing MLIP to include a new chemical element requires generating thousands of new reference calculations and retraining the entire model — a process that represents a significant computational bottleneck.
@@
Foundational MLIPs resolve this problem by embedding complex chemistry into a low-dimensional space; a long-standing concept in materials simulation. Early tight-binding models, for example, used the valence electron count to effectively describe chemical trends and structural stability across multiple elements Pettifor (1970). Modern foundational MLIPs build upon this legacy, using multi-dimensional embeddings to leverage the inherent correlations between elements. This approach is remarkably effective, enabling the entire periodic table’s chemistry to be captured in few dimensions.Pettifor (1986, 1995); Seiser et al. (2011); Bialon et al. (2016); Faber et al. (2015); Parsaeifard et al. (2021); Lopanitsyna et al. (2023a); Cerqueira et al. (2025)
@@
The first universal force field was published more than 30 years agoRappé et al. (1992). Parameterizations of MLIPs across the periodic table started to appear when large training datasets became available Chen and Ong (2022); Deng et al. (2023a); Batatia et al. (2023); Barroso-Luque et al. (2024a); Yang et al. (2024); Kim et al. (2024); Yin et al. (2025); Zhang et al. (2025); Fu et al. (2025); Mazitov et al. (2025); Liang et al. (2025).
@@
To date the development of universal MLIPs was tightly bound to the progress in message passing graph neural networks Gasteiger et al. (2020); Anderson et al. (2019); Lubbers et al. (2018); Thomas et al. (2018); Batzner et al. (2022); Satorras et al. (2021); Unke and Meuwly (2019); Schütt et al. (2017); Haghighatlari et al. (2022); Schütt et al. (2021); Batzner et al. (2022); Gasteiger et al. (2022); Chmiela et al. (2017); Pozdnyakov and Ceriotti (2023); Nigam et al. (2022); Batatia et al. (2025a); Bochkarev et al. (2022a, 2024), despite the fact that universal parameterizations across the periodic table are in principle independent of MLIP architecture. The first universal MLIPs that built on the Atomic Cluster Expansion (ACE) Drautz (2019) became available two years ago, within the framework of Multi-ACE Batatia et al. (2025b) that employs general many-body messages and as implemented in MACE Batatia et al. (2022, 2023).
@@
Here we employ the Graph Atomic Cluster Expansion (GRACE)Bochkarev et al. (2024). By extending ACE to tree-graphs, GRACE stands out from graph neural networks by providing a complete basis for the parameterization of atomic interactions as a function of atomic positions and chemical species. By straightforward tensor decomposition of the GRACE expansion coefficients one directly obtains sparse representations with efficient chemical embedding that can be evaluated recursively. The recursive evaluation of graph basis functions can be understood as message passing and because of its complete basis, GRACE is able to rationalize and represent other message passing graph neural networks architectures in general. As GRACE further facilitates recursive evaluation of effective ACE on each message passing layer, it benefits from linear scaling with the number of recursion layers as well as linear scaling with the complexity of ACE messages within each layer for efficient double-recursive evaluation.
@@
Training data for universal force fields needs to cover the periodic table comprehensively. Only few publicly available datasets are suitable, notably the Materials Project Jain et al. (2013); Deng et al. (2023b), Alexandria Schmidt et al. (2023); Wang et al. (2023) and the OMat24 Barroso-Luque et al. (2024b) datasets that are in the focus of our work here, but also the Open Quantum Mechanical Database Kirklin et al. (2015) (OQMD) and AFLOWLIB Curtarolo et al. (2012) and more recent additions such as MatPES Kaplan et al. (2025) and MP-ALOE Kuner et al. (2025).
@@
In this work we present a number of GRACE models with varying complexity that were parameterized on the OMat24 Barroso-Luque et al. (2024b), Alexandria Schmidt et al. (2023); Wang et al. (2023) and MPTraj Deng et al. (2023b) datasets and can serve as a foundation models for atomistic modelling in materials science.
@@ II.1 Foundational GRACE interatomic potentials
Developing foundational MLIPs capable of accurate predictions across a wide array of chemical elements and diverse structures requires exceptionally large and varied datasets. The GRACE framework is designed to effectively manage this inherent complexity. Our primary training source was the OMat24 dataset Barroso-Luque et al. (2024b), which currently is the largest publicly available compilation for materials property prediction. It encompasses 110 million DFT calculations, primarily computed with VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b) with the GGA-PBE functionalPerdew et al. (1996), including Hubbard U corrections for specific oxides and fluorides, consistent with Materials Project defaults Jain et al. (2013). Importantly, OMat24 extends beyond near-equilibrium structures, distinguishing it from datasets like Alexandria and MPTraj, but covers same 89 elements as those two. The dataset’s diversity comes from its generation methods, which include Boltzmann sampling of structures with randomly displaced atomic positions, ab initio molecular dynamics (AIMD), and subsequent relaxations of these configurations.
@@
While OMat24 provides a robust foundation, its DFT and pseudopotential settings differ from those used in Alexandria and the Materials Project. To address this, we fine-tuned the OMAT-base models using a combined dataset of MPTraj and a subsampled Alexandria (sAlex) dataset Schmidt et al. (2023); Wang et al. (2023); Barroso-Luque et al. (2024b). The sAlex subset was curated to prevent data leakage with the WBM test set Wang et al. (2021), a crucial step to ensure model compatibility with Matbench Discovery Riebesell et al. (2025). The resulting models are denoted with suffix “-OAM”. These GRACE models are designed to serve as robust foundational interatomic potentials and to provide uniform accuracy across a broad range of chemical compositions and structural configurations. More details on the models architecture and parameterization can be found in Supplementary Information IX.1.
@@ II.2 Validation
Foundational MLIPs must demonstrate uniform accuracy across multiple application domains. In this section, we present several critical validation tests. We evaluate our models against the MatBench Discovery benchmark Riebesell et al. (2025) for formation energies and thermodynamic stability and the $\kappa$-SRME Póta et al. (2024) test for thermal conductivity, which reflects second- and third-order derivatives. We further determine the performance of our models on elastic properties and for non-equilibrium and defective configurations by predicting formation energies of grain boundaries, surfaces, and point defects in pure elements. This suite of tests provides a good assessment of the models’ capabilities and limitations.
@@ II.2.1 MatBench Discovery benchmark
Matbench Discovery Riebesell et al. (2025) serves as a benchmark for high-throughput discovery of stable inorganic crystals. It is specifically designed to evaluate the efficacy of various foundational MLIPs in predicting formation energy and thermodynamic stability of novel crystal structures. The benchmark task involves geometry optimization of structures sourced from the WBM dataset, a diverse collection of 257,000 candidate crystal structures spanning a wide range of compositions. The central goal is to predict formation energies and assess the stability of these structures relative to the original convex hull from Materials Project Jain et al. (2013) with high fidelity. Achieving this requires models to accurately predict potential energy, atomic forces, and stress tensors, with an accuracy comparable to the Materials Project’s density functional theory (DFT) calculations.
Figure 1 (left panel) illustrates the performance of GRACE models, specifically the OAM-fine-tuned versions, against other publicly available foundational MLIPs in relation to their computational efficiency. The figure shows that GRACE models consistently occupy the Pareto optimality front, demonstrating a superior balance of performance and computational speed. This efficiency can be further enhanced by using the LAMMPS molecular dynamics code Thompson et al. (2022), with specific timings available in Table 1. While the F1 metric reflects a model’s accuracy for identifying stable structures, the mean absolute error (MAE) of formation energies offers a more general measure of accuracy. The performance of the models for formation energy MAE, which shows a similar trend with GRACE models on the Pareto front, is provided in the supplementary materials.
@@ II.2.2 Thermal conductivity
To evaluate the ability of foundational MLIPs to predict force-dependent properties like phonons and anharmonic thermal conductivity, we used the symmetric relative mean error $\kappa_{\mathrm{SRME}}$ metric Póta et al. (2024). This test quantifies a model’s performance by predicting the thermal conductivity $\kappa$ across 103 binary structures. The thermal conductivity values are calculated from forces predicted by the foundational MLIPs and subsequently analyzed using the phono3py software Togo et al. (2015, 2023). Accurate thermal conductivity predictions serve as a strong indicator of a model’s performance for other simulation tasks, such as modeling metal-organic frameworks Kraß et al. (2025).
@@ II.2.3 Elastic properties
Predicting elastic moduli is a crucial validation test for interatomic potentials. We categorize elastic constants into three subgroups: longitudinal ($C_{11},C_{22},C_{33}$), Poisson’s ratio-related ($C_{12},C_{13},C_{23}$), and shear ($C_{44},C_{55},C_{66}$). Because these groups often have varying magnitudes, we primarily focused on the symmetric relative mean error (SRME) and MAE within each subgroup. Figure 2 presents both metrics with respect to reference data from the Materials Project De Jong et al. (2015).
@@ II.2.4 Grain boundaries
To assess the performance of foundational MLIPs for bulk structural defects, we utilized an existing dataset of grain boundary formation energies ($\gamma_{\mathrm{GB}}$ computed for pure metals Zheng et al. (2020). The models’ accuracy was quantified by calculating both the $\gamma_{\mathrm{GB}}$-SRME and the mean absolute error $\Delta\gamma_{\mathrm{GB}}$. The results are presented in Fig. 3.
@@ II.2.5 Surface energies
We assessed the ability of foundational MLIPs to predict open structures, such as surfaces, using surface energies ($\gamma_{\mathrm{surf}}$) for pure elements Tran et al. (2016). Model accuracy was quantified by calculating both the symmetric relative mean error ($\gamma_{\mathrm{surf}}$-SRME) and the mean absolute error ($\Delta\gamma_{\mathrm{surf}}$), with results presented in Fig. 4. Here, the relative error $\gamma_{\mathrm{surf}}$-SRME ranged from 0.168 for the ORBv3-cons-inf-mpa model to 0.279 for the MACE-OMAT-0 model. The absolute error $\Delta\gamma_{\mathrm{surf}}$ typically varied from 8 to 14 meV/Å2, with a few exceptions noted for MACE-MPA-0, MatterSim, and eqV2_M. A consistent finding across all models was poor $\gamma_{\mathrm{surf}}$-SRME metrics for Potassium (K), Rubidium (Rb), Cesium (Cs), and Indium (In). More detailed information can be found in the Supplementary Information.
@@ II.2.6 Point defects
We used systematically computed formation energies for self-interstitials (SIA) and vacancies in BCC Ma and Dudarev (2019) and FCC Ma and Dudarev (2021) metals available in the literature as references. Given the varying scales of these formation energies across different defect types and metals, we used the symmetric relative mean error metrics, ESIA-SRME and Evac-SRME, as main measure of accuracy.
@@ II.3 Long-time stability of MD
Long-time stability in molecular dynamics (MD) simulations is essential for accurately capturing the dynamical behavior and thermodynamic properties that cannot be fully assessed by static property benchmarks or short simulations. To demonstrate the out-of-the-box MD stability and performance of the GRACE foundational MLIPs, we conducted a 1 ns MD simulation of a FLiBe cell containing approximately three thousands atoms at 973 K, utilizing the GRACE-2L-OMAT-L model. We compared the resulting radial distribution functions (RDFs) to reference AIMD data from Ref. Lam et al. (2021); Rodriguez et al. (2021), as shown in Fig. 6. Additionally, we estimated diffusion coefficients for each element from our MD simulation, obtaining values of 1.33, 1.58, and $5.86\times 10^{-5}$ cm2s-1 for Be, F, and Li, respectively. These values align well with the AIMD results ($0.83\pm 0.1$, $1.73\pm 0.17$ and $5.67\pm 0.52\times 10^{-5}$ cm2s-1, respectively Lam et al. (2021)). This simulation confirms the model’s stability and its accuracy in predicting both structural and dynamical properties over extended timescales.
Figure 6: RDF of FLiBe salt from MD at 973K using GRACE in comparison to AIMD-DFT results Lam et al. (2021); Rodriguez et al. (2021).
@@ II.4 Computational performance
We also measured the computational performance of other foundational MLIPs by running ten MD steps in ASE Larsen et al. (2017) of W-BCC supercells with different numbers of atoms. These tests utilized different GPU hardware, including the commodity RTX3060 (12 GB), and L40s (40 GB), A100 (80 GB), and H200 (141 GB). The resulting execution times are presented in Table 2. GRACE models deliver excellent performance even on commodity GPUs like the RTX 3060, despite consistently operating with FP64 (double-precision) accuracy.
@@ II.5 Fine-tuning
Here we demonstrate the fine-tuning of GRACE-2L foundational MLIP for the Al-Li binary system Menon et al. (2024) and a hydrogen combustion dataset Guan et al. (2022).

### v2

@@ I Introduction
The ability to predict materials properties from atomistic simulations is essential for modern materials design. Machine learning interatomic potentials (MLIPs), trained on data from electronic structure methods like density functional theory, have recently emerged as a powerful tool, achieving excellent accuracy for diverse systems Behler07; Bartok2010; Shapeev16; Drautz19; unkeMachineLearningForce2021; musaelianLearningLocalEquivariant2023; jacobsPracticalGuideMachine2025; kulichenkoDataGenerationMachine2024; WANG2024109673; Thiemann_2025; yuan2025foundationmodelsatomisticsimulation. However, the applicability of conventional MLIPs is constrained by their limited elemental scope. Extending an existing MLIP to include a new chemical element requires generating thousands of new reference calculations and retraining the entire model - a process that represents a significant computational bottleneck.
@@
Foundational MLIPs resolve this problem by embedding complex chemistry into a low-dimensional space; a long-standing concept in materials simulation. Early tight-binding models, for example, used the valence electron count to effectively describe chemical trends and structural stability across multiple elements pettiforTheoryCrystalStructures1970. Modern foundational MLIPs build upon this legacy, using multi-dimensional embeddings to leverage the inherent correlations between elements. This approach is remarkably effective, enabling the entire periodic table’s chemistry to be captured in few dimensions.DGPettifor_1986; pettiforbook; SEISER2011749; acs.chemmater.5b04299; qua.24917; Parsaeifard_2021; PhysRevMaterials.7.045802; cerqueira2025nonorthogonalrepresentationchemicalspace
@@
The first universal force field was published more than 30 years agorappeUFFFullPeriodic1992. Parameterizations of MLIPs across the periodic table started to appear when large training datasets became available chen2022universal; dengCHGNetPretrainedUniversal2023; batatiaFoundationModelAtomistic2024; barroso2024open; yang2024mattersim; kim2024data; yin2025alphanet; zhang2025graph; fu2025learning; mazitovPETMADLightweightUniversal2025; liang2025nep89.
@@
To date the development of universal MLIPs was tightly bound to the progress in message passing graph neural networks Klicpera2020DimeNet; Anderson2019Cormorant; Lubbers2018; Thomas2018TensorField; Batzner2022nequip; Satorras2021EGNN; Unke2019Physnet; Schutt2017schnet; Haghighatlari2021newtonnet; Schutt2021Painn; Batzner2022nequip; Klicpera2022gemnet; Chmiela2017machine; Pozdnyakov2023smooth; Nigam2022unified; Batatia2022design; bochkarev2022multilayer; bochkarev2024, despite the fact that universal parameterizations across the periodic table are in principle independent of MLIP architecture. The first universal MLIPs that built on the Atomic Cluster Expansion (ACE) Drautz19 became available two years ago, within the framework of Multi-ACE batatiaDesignSpaceE3equivariant2025 that employs general many-body messages and as implemented in MACE batatia2022mace; batatiaFoundationModelAtomistic2024.
@@
Here we employ the Graph Atomic Cluster Expansion (GRACE)bochkarev2024. By extending ACE to tree-graphs, GRACE stands out from graph neural networks by providing a complete basis for the parameterization of atomic interactions as a function of atomic positions and chemical species. By straightforward tensor decomposition of the GRACE expansion coefficients one directly obtains sparse representations with efficient chemical embedding that can be evaluated recursively. The recursive evaluation of graph basis functions can be understood as message passing and because of its complete basis, GRACE is able to rationalize and represent other message passing graph neural networks architectures in general. As GRACE further facilitates recursive evaluation of effective ACE on each message passing layer, it benefits from linear scaling with the number of recursion layers as well as linear scaling with the complexity of ACE messages within each layer for efficient double-recursive evaluation.
@@
Training data for universal force fields needs to cover the periodic table comprehensively. Only few publicly available datasets are suitable, notably the Materials Project jain2013commentary; deng2023chgnet, Alexandria schmidt2023machine; wang2023symmetry and the OMat24 barroso_omat24 datasets that are in the focus of our work here, but also the Open Quantum Mechanical Database kirklin2015open (OQMD) and AFLOWLIB curtarolo2012aflowlib and more recent additions such as MatPES MATPES and MP-ALOE kuner2025mpaloe.
@@
In this work we present a number of GRACE models with varying complexity that were parameterized on the OMat24 barroso_omat24, Alexandria schmidt2023machine; wang2023symmetry and MPTraj deng2023chgnet datasets and can serve as a foundation models for atomistic modelling in materials science.
@@ II.1 Foundational GRACE interatomic potentials
Developing foundational MLIPs capable of accurate predictions across a wide array of chemical elements and diverse structures requires exceptionally large and varied datasets. The GRACE framework is designed to effectively manage this inherent complexity. Our primary training source was the OMat24 dataset barroso_omat24, which currently is the largest publicly available compilation for materials property prediction. It encompasses 110 million DFT calculations, primarily computed with VASP kressabinitio1993; kresse1996efficiency; kresse1996efficient with the GGA-PBE functionalPerdew1996, including Hubbard U corrections for specific oxides and fluorides, consistent with Materials Project defaults jain2013commentary. Importantly, OMat24 extends beyond near-equilibrium structures, distinguishing it from datasets like Alexandria and MPTraj, but covers same 89 elements as those two. The dataset’s diversity comes from its generation methods, which include Boltzmann sampling of structures with randomly displaced atomic positions, ab initio molecular dynamics (AIMD), and subsequent relaxations of these configurations.
@@
While OMat24 provides a robust foundation, its DFT and pseudopotential settings differ from those used in Alexandria and the Materials Project. To address this, we fine-tuned the OMAT-base models using a combined dataset of MPTraj and a subsampled Alexandria (sAlex) dataset schmidt2023machine; wang2023symmetry; barroso_omat24. The sAlex subset was curated to prevent data leakage with the WBM test set wang2021predicting, a crucial step to ensure model compatibility with Matbench Discovery riebesell2025framework. The resulting models are denoted with suffix “-OAM”. These GRACE models are designed to serve as robust foundational interatomic potentials and to provide uniform accuracy across a broad range of chemical compositions and structural configurations. More details on the models architecture and parameterization can be found in Supplementary Information X.1.
@@ II.2 Validation
Foundational MLIPs must demonstrate uniform accuracy across multiple application domains. In this section, we present several critical validation tests. We evaluate our models against the MatBench Discovery benchmark riebesell2025framework for formation energies and thermodynamic stability and the $\kappa$-SRME pota2024thermal test for thermal conductivity, which reflects second- and third-order derivatives. We further determine the performance of our models on elastic properties and for non-equilibrium and defective configurations by predicting formation energies of grain boundaries, surfaces, and point defects in pure elements. This suite of tests provides a good assessment of the models’ capabilities and limitations.
@@ II.2.1 MatBench Discovery benchmark
Matbench Discovery riebesell2025framework serves as a benchmark for high-throughput discovery of stable inorganic crystals. It is specifically designed to evaluate the efficacy of various foundational MLIPs in predicting formation energy and thermodynamic stability of novel crystal structures. The benchmark task involves geometry optimization of structures sourced from the WBM dataset, a diverse collection of 257,000 candidate crystal structures spanning a wide range of compositions. The central goal is to predict formation energies and assess the stability of these structures relative to the original convex hull from Materials Project jain2013commentary with high fidelity. Achieving this requires models to accurately predict potential energy, atomic forces, and stress tensors, with an accuracy comparable to the Materials Project’s density functional theory (DFT) calculations.
Figure 1 (left panel) illustrates the performance of GRACE models, specifically the OAM-fine-tuned versions, against other publicly available foundational MLIPs in relation to their computational efficiency. The figure shows that GRACE models consistently occupy the Pareto optimality front, demonstrating a superior balance of performance and computational speed. This efficiency can be further enhanced by using the LAMMPS molecular dynamics code LAMMPS, with specific timings available in Table 1. While the F1 metric reflects a model’s accuracy for identifying stable structures, the mean absolute error (MAE) of formation energies offers a more general measure of accuracy. The performance of the models for formation energy MAE, which shows a similar trend with GRACE models on the Pareto front, is provided in the supplementary materials.
@@ II.2.2 Thermal conductivity
To evaluate the ability of foundational MLIPs to predict force-dependent properties like phonons and anharmonic thermal conductivity, we used the symmetric relative mean error $\kappa_{\mathrm{SRME}}$ metric pota2024thermal. This test quantifies a model’s performance by predicting the thermal conductivity $\kappa$ across 103 binary structures. The thermal conductivity values are calculated from forces predicted by the foundational MLIPs and subsequently analyzed using the phono3py software togo2015distributions; togo2023implementation. Accurate thermal conductivity predictions serve as a strong indicator of a model’s performance for other simulation tasks, such as modeling metal-organic frameworks krass2025mofsimbench.
@@ II.2.3 Elastic properties
Predicting elastic moduli is a crucial validation test for interatomic potentials. We categorize elastic constants into three subgroups: longitudinal ($C_{11},C_{22},C_{33}$), Poisson’s ratio-related ($C_{12},C_{13},C_{23}$), and shear ($C_{44},C_{55},C_{66}$). Because these groups often have varying magnitudes, we primarily focused on the symmetric relative mean error (SRME) and MAE within each subgroup. Figure 2 presents both metrics with respect to reference data from the Materials Project de2015charting.
@@ II.2.4 Grain boundaries
To assess the performance of foundational MLIPs for bulk structural defects, we utilized an existing dataset of grain boundary formation energies ($\gamma_{\mathrm{GB}}$ computed for pure metals zheng2020grain. The models’ accuracy was quantified by calculating both the $\gamma_{\mathrm{GB}}$-SRME and the mean absolute error $\Delta\gamma_{\mathrm{GB}}$. The results are presented in Fig. 3.
@@ II.2.5 Surface energies
We assessed the ability of foundational MLIPs to predict open structures, such as surfaces, using surface energies ($\gamma_{\mathrm{surf}}$) for pure elements tran2016surface. Model accuracy was quantified by calculating both the symmetric relative mean error ($\gamma_{\mathrm{surf}}$-SRME) and the mean absolute error ($\Delta\gamma_{\mathrm{surf}}$), with results presented in Fig. 4. Here, the relative error $\gamma_{\mathrm{surf}}$-SRME ranged from 0.168 for the ORBv3-cons-inf-mpa model to 0.279 for the MACE-OMAT-0 model. The absolute error $\Delta\gamma_{\mathrm{surf}}$ typically varied from 8 to 14 meV/Å2, with a few exceptions noted for MACE-MPA-0, MatterSim, and eqV2_M. A consistent finding across all models was poor $\gamma_{\mathrm{surf}}$-SRME metrics for Potassium (K), Rubidium (Rb), Cesium (Cs), and Indium (In). More detailed information can be found in the Supplementary Information.
@@ II.2.6 Point defects
We used systematically computed formation energies for self-interstitials (SIA) and vacancies in BCC ma2019universality and FCC ma2021nonuniversal metals available in the literature as references. Given the varying scales of these formation energies across different defect types and metals, we used the symmetric relative mean error metrics, ESIA-SRME and Evac-SRME, as main measure of accuracy.
@@ II.3 Long-time stability of MD
[…]
Figure 6: RDF of FLiBe salt from MD at 973K using GRACE in comparison to AIMD-DFT results lam2021modeling; rodriguez2021thermodynamic.
@@ II.4 Computational performance
We also measured the computational performance of other foundational MLIPs by running ten MD steps in ASE ase-paper of W-BCC supercells with different numbers of atoms. These tests utilized different GPU hardware, including the commodity RTX3060 (12 GB), and L40s (40 GB), A100 (80 GB), and H200 (141 GB). The resulting execution times are presented in Table 2. GRACE models deliver excellent performance even on commodity GPUs like the RTX 3060, despite consistently operating with FP64 (double-precision) accuracy. As shown in Fig.1, the use of ASE can introduce substantial computational overhead, stemming primarily from Python’s execution speed and suboptimal neighbor list construction algorithms.
@@ II.5 Fine-tuning
Here we demonstrate the fine-tuning of GRACE-2L foundational MLIP for the Al-Li binary system menon2024electrons and a hydrogen combustion dataset guan2022benchmark.